# Breaking Imitation Bottlenecks: Reinforced Diffusion Powers Diverse Trajectory Generation

### v2

@@ 4.3.2 GRPO for Trajectory Planning
Despite their strong capability in multi-mode trajectories generation, diffusion models still face two critical limitations in end-to-end autonomous driving (E2E-AD): (i) mode collapse, where generated trajectories converge to similar patterns, and (ii) lack of safety awareness, leading to kinematically feasible yet unsafe motions. These issues largely stem from reliance on single-GT supervision and the absence of explicit guidance for task-specific objectives such as behavioral diversity and safety.
@@
Diversity Reward $r_{\text{div}}$. To encourage multi-mode trajectories generation and mitigate mode collapse, we design a diversity reward $r_{\text{div}}$ that quantifies the pairwise dissimilarity among predicted trajectories. Given a set of $M$ predicted trajectories $\{\hat{\tau}^{(1)},\hat{\tau}^{(2)},\dots,\hat{\tau}^{(M)}\}$, we compute:
@@
Safety Reward $r_{\text{safe}}$. To ensure the feasibility and risk-awareness of generated trajectories, we introduce a safety reward that penalizes predicted paths with low clearance from static or dynamic obstacles. We define a differentiable distance-based cost function using a safety map $D_{\text{safe}}(\mathbf{x})$, which encodes inverse proximity to obstacles. For a trajectory $\hat{\tau}={\hat{x}*t}*{t=1}^{T}$, the safety reward is computed as:
$r_{\text{safe}}(\hat{\tau})=-\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}\left[D_{\text{safe}}(\hat{x}*t)<d*{\text{thresh}}\right],$ (19)
@@
Planing Loss for Diverse Trajectories. The total reward for a predicted trajectory $\hat{\tau}^{(m)}$ is defined as:
$r(\hat{\tau}^{(m)})=r_{\text{div}}(\hat{\tau}^{(m)})+\lambda_{\text{safe}}\cdot r_{\text{safe}}(\hat{\tau}^{(m)}).$ (20)
The reward is integrated into a reinforcement learning loss under the GRPO framework:
$\mathcal{L}_{\text{RL}}=-\mathbb{E}_{\hat{\tau}^{(m)}\sim\pi_{\theta}}\left[\text{GRPO}(\hat{\tau}^{(m)},r(\hat{\tau}^{(m)}))\right]$ (21)
This formulation enables the diffusion model to adaptively steer its sampling process toward trajectories that are diverse, feasible, and physically plausible.
@@
$\mathcal{L}_{\text{total}}=\lambda_{\text{match}}\mathcal{L}_{\text{match}}+\lambda_{\text{RL}}\mathcal{L}_{\text{RL}}$ (22)
Here, $\lambda_{\text{match}}$ and $\lambda_{\text{RL}}$ are weighting coefficients balancing the contributions of trajectory matching and policy optimization.
Our reinforcement-augmented training framework addresses key limitations of diffusion-based planners in E2E-AD. Through reward-guided GRPO optimization, we impose safety and diversity constraints that are not captured by conventional MSE-based training. This hybrid supervision strategy bridges the gap between behavior cloning and policy-level optimization, leading to a more robust and deployable planning solution.
@@ 5.2 Evaluation Metrics
Overall, a higher Diversity value indicates greater dispersion among the predicted trajectories, reflecting a richer coverage of intent modes such as lane changes, braking, and yielding. When Diversity approaches 1, the predicted modes are highly diverse and well-separated in the trajectory space. Conversely, a value near 0 suggests that the predicted trajectories are highly similar, indicating a lack of behavioral diversity. Notably, the Diversity Metric is only applicable to multi-mode trajectories E2E-AD methods.
@@ 5.3 Implementation Details
TABLE I: $\operatorname{\textbf{Open-Loop}}$, $\operatorname{\textbf{Closed-Loop}}$ results and $\operatorname{Multi-Ability}$ results on $\operatorname{Bench2Drive}$ (V0.0.3) under base training set. ‘mmt’ refers multi-mode trajectories variant of $\operatorname{VAD}$ and denotes the re-implementation. * denotes expert feature distillation. ‘DS’ denotes Driving Score. ‘SR’ denotes Success Rate. ‘Effi’ denotes Efficiency. ‘Comf’ denotes Comfortness. ‘Merg.’ denotes Merging. ‘Overta.’ denotes Overtaking. ‘Emerge.’ denotes Emergency Brake. It is worth noting that the Diversity Metric ($\operatorname{\textit{Div.}^{(t)}}\uparrow$) is applicable only to multi-mode E2E-AD methods.
@@
WoTE*[64] ST IL Arxiv 2025 - - 61.71 31.36 - - - - - - - -
DiffAD*[39] MT IL Arxiv 2025 1.55 - 67.92 38.64 - - - - - - - -
@@
TABLE II: Comparison on planning-oriented $\operatorname{NAVSIM}$ navtest split with $\operatorname{\textbf{Closed-Loop}}$ metrics. ‘mmt’ refers multi-mode trajectories variant of $\operatorname{TransFuser}$ and denotes the re-implementation.
@@
TABLE III: Planning results on the $\operatorname{\textbf{Open-Loop}}$ $\operatorname{NuScenes}$ [51] validation dataset. It is worth noting that the Diversity Metric ($\operatorname{\textit{Div.}^{(t)}}\uparrow$) is applicable only to multi-mode E2E-AD methods like MomAD [58], DiffusionDrive [4], and SparseDrive [3].
@@
TABLE IX: Ablation Study on Policy-Aware Diffusion Generator on NAVSIM dataset. ‘Condition’ denotes map and agent conditioning in the diffusion model, while ‘Reference GTs’ refers to reference trajectory guidance.
@@
+ Condition 96.2 96.3 90.8 100 81.2 86.3
+ Reference GTs 97.6 95.7 91.5 100 81.6 86.4
$\operatorname{DIVER}$ 98.5 96.5 94.9 100 82.6 88.3
@@
TABLE XI: Ablation Study on Loss Function Design on NuScenes dataset.
@@
TABLE XII: Ablation Study on Loss Function Design on NAVSIM dataset.
@@ 5.4 Main Results
To comprehensively evaluate the multi-mode trajectories generation capability of our proposed DIVER, we conduct closed-loop experiments on two representative E2E-AD benchmarks: Bench2Drive and NAVSIM, as shown in Tables I and II. In Table III, we further present open-loop results on the NuScenes dataset. Additionally, Table I includes supplementary open-loop evaluation results on Bench2Drive. It is important to note that we do not adopt the conventional L2 distance in open-loop evaluation. This is because L2 distance merely reflects the proximity between a predicted trajectory and a single GT trajectory, failing to capture the diversity and spread of multi-mode trajectories predictions. Relying solely on this metric compromises the core objective of multi-mode trajectory generation—namely, to cover a wide range of plausible future intents.
@@ 5.4.1 Close-Loop Results
NAVSIM (Close-Loop). As shown in Table II, our proposed DIVER achieves the better overall performance across all closed-loop metrics on the NAVSIM planning-oriented benchmark. This benchmark evaluates planners under non-reactive simulations using multiple dimensions, including No at-fault Collisions (NC), Drivable Area Compliance (DAC), Time-to-Collision (TTC), Comfort (Comf.), Ego Progress (EP), and the Planning-Driven Metric Score (PDMS) [55]. DIVER attains the highest PDMS score of 88.3, outperforming strong baselines such as $\text{TransFuser}_{\text{mmt}}$ and DiffusionDrive, which achieve 87.5 and 86.1, respectively. In terms of safety, DIVER leads with the highest NC score of 98.5, indicating fewer at-fault collisions. Moreover, DIVER demonstrates excellent planning feasibility and smoothness with top scores in DAC (96.5) and Comf. (100), while maintaining strong forward progress (EP: 82.6). These results highlight DIVER’s ability to balance safety, comfort, and goal-directed planning, demonstrating its effectiveness in realistic driving simulations. The improvements across all aspects suggest that our method not only generates diverse multi-mode trajectories but also ensures their practical executability and robustness in closed-loop scenarios.
@@ 5.6.1 Ablation Study on PADG
Impact of the Number of Reference GTs. As shown in Figure 5, increasing the number of Reference GTs leads to consistent gains in Driving Score, Success Rate, and complex maneuvers (e.g., merging, overtaking). Performance stabilizes around 6–7 GTs, indicating that moderate reference diversity effectively enhances planning robustness and behavioral richness.
@@ 5.6.2 Ablation Study on Loss Function Design
As shown in Tables XI and XII, loss function design has a significant impact on model performance. On the NuScenes dataset, the standard L1 loss fails to exploit the potential of diffusion and reference GTs, yielding limited diversity (Avg. Div.↑ = 0.14). Even when combined with RL (PPO), improvements are marginal. In contrast, replacing L1 with the proposed matching loss $\mathcal{L}_{\text{match}}$ enables the model to benefit from trajectory diversity, yielding higher $\operatorname{\textit{Div.}^{(t)}}$ (up to 0.21). Further integrating RL with diversity and safety rewards enhances performance, especially using GRPO, which outperforms PPO in both diversity and planning quality. On NAVSIM, $\mathcal{L}_{\text{match}}+\mathcal{L}_{\text{RL(GRPO)}}$ achieves the best PDMS (88.3), confirming that this design effectively balances realism, diversity, and safety.

### v3

@@ 4.3.2 GRPO for Trajectory Planning
Despite their strong capability in multi-mode trajectory generation, diffusion models still face two critical limitations in end-to-end autonomous driving (E2E-AD): (i) mode collapse, where generated trajectories converge to similar patterns, and (ii) lack of safety awareness, leading to kinematically feasible yet unsafe motions. These issues largely stem from reliance on single-GT supervision and the absence of explicit guidance for task-specific objectives such as behavioral diversity and safety.
@@
Hydra-MDP [11] promotes multi-mode trajectory generation without relying on imitation learning by directly distilling PDMS-based metrics into the policy network, where such metric distillation can be regarded as a single-step reinforcement learning process. Following the principles of Hydra-MDP [11] and GTRS [57], we incorporate a PDMS-guided reinforcement optimization into our framework, implemented on the NAVSIM dataset. Specifically, we reuse the PDMS from Hydra-MDP to compute driving-quality metrics including no at-fault collisions (NC), drivable area compliance (DAC), time-to-collision (TTC), comfort (Comf.) and ego progress (EP), and provide online reinforcement guidance within the GRPO optimization loop. Unlike Hydra-MDP, which distills these metrics into a static policy, our approach dynamically integrates the PDMS-based rewards during training, enabling the diffusion policy to adaptively align with high-level safety and driving-quality objectives. This PDMS-guided reinforcement mechanism complements the trajectory-level rewards of GRPO, thereby enhancing multi-mode diversity, safety awareness, and behavioral consistency.
@@
Diversity Reward $r_{\text{div}}$. To encourage multi-mode trajectory generation and mitigate mode collapse, we design a diversity reward $r_{\text{div}}$ that quantifies the pairwise dissimilarity among predicted trajectories. Given a set of $M$ predicted trajectories $\{\hat{\tau}^{(1)},\hat{\tau}^{(2)},\dots,\hat{\tau}^{(M)}\}$, we compute:
@@
Safety Reward $r_{\text{safe}}$. To ensure the feasibility and risk-awareness of generated trajectories, we introduce a safety reward that penalizes predicted paths with low clearance from static or dynamic obstacles. We define a differentiable distance-based cost function using a safety map $D_{\text{safe}}(\mathbf{x})$, which encodes inverse proximity to obstacles. For a trajectory $\hat{\tau}={\hat{x}\cdot t}\cdot{t=1}^{T}$, the safety reward is computed as:
$r_{\text{safe}}(\hat{\tau})=-\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}\left[D_{\text{safe}}(\hat{x}\cdot t)<d_{\text{thresh}}\right],$ (16)
@@
Trajectory Consistency Reward $r_{\text{TC}}$. To maintain temporal coherence across consecutive planning cycles, we introduce a trajectory consistency reward that enforces alignment between the current predicted trajectory $\hat{\tau}_{t}$ and the previous executed (or predicted) trajectory $\hat{\tau}_{t-1}$. Sudden directional shifts between adjacent planning steps may lead to unstable or oscillatory driving behaviors. The reward is formulated as:
$r_{\text{TC}}=-\frac{1}{T}\sum_{k=1}^{T}\left\|\hat{x}_{t,k}-\hat{x}_{t-1,k}\right\|_{2},$ (17)
where $\hat{x}_{t,k}$ denotes the $k$-th waypoint of the trajectory at time $t$. By minimizing inter-frame deviations, this reward encourages smooth transitions and consistent motion patterns over time, reducing unnecessary steering oscillations and improving temporal stability.
Lane Keeping Reward $r_{\text{LK}}$. To further constrain the ego vehicle to follow reasonable lane-keeping behaviors, we design a lane keeping reward that penalizes frequent or unnecessary lane changes. Let $\mathcal{L}(\hat{x}_{t}^{i})$ denote the distance from the $i$-th predicted waypoint $\hat{x}_{t}^{i}$ to the lane centerline. And $\tau_{LK}$ denotes the distance threshold for determining lane-keeping compliance, which is set to 0.5. We apply the following rule to each waypoint $\hat{x}_{t}^{i}$ to obtain a binary (0/1) array arrarr.
$arr=[((\mathcal{L}(\hat{x}_{t}^{i})-\tau_{LK})>0),...],\ i=0,1,...,k,$ (18)
We iterate through the array arrarr and compute the maximum $l_{arr}$ length of consecutive ones in the array. Then $r_{\text{LK}}$ can be computed by:
$r_{\text{LK}}=1-((l_{arr}-\delta_{LK})>0),$ (19)
where $\delta_{LK}$ represents the threshold and is set to 4.
Planning Loss for Diverse and Consistent Trajectories. The total reward for a predicted trajectory $\hat{\tau}^{(m)}$ is defined as:
$r(\hat{\tau}^{(m)})=\;$ $\lambda_{\text{div}}\cdot r_{\text{div}}(\hat{\tau}^{(m)})+\lambda_{\text{safe}}\cdot r_{\text{safe}}(\hat{\tau}^{(m)})$
$+\lambda_{\text{TC}}\cdot r_{\text{TC}}(\hat{\tau}^{(m)})+\lambda_{\text{LK}}\cdot r_{\text{LK}}(\hat{\tau}^{(m)}),$ (20)
where $r_{\text{div}}$, $r_{\text{safe}}$, $r_{\text{TC}}$, and $r_{\text{LK}}$ denote the diversity, safety, trajectory consistency, and lane-keeping rewards, respectively. It is worth noting that DIVER is primarily designed to mitigate trajectory mode collapse and enhance trajectory diversity. Accordingly, DIVER incorporates the reward terms $r_{\text{TC}}$ and $r_{\text{LK}}$. In the ablation study, we further compare the effects of $r_{\text{TC}}$, $r_{\text{LK}}$, and PDMS-guided strategies. Overall, our DIVER framework provides a scalable and extensible reward design, which can be flexibly adapted or revised to align with different objectives and human values.
The total reward is integrated into a reinforcement learning loss under the GRPO framework:
$\mathcal{L}_{\text{RL}}=-\mathbb{E}_{\hat{\tau}^{(m)}\sim\pi_{\theta}}\left[\text{GRPO}\big(\hat{\tau}^{(m)},r(\hat{\tau}^{(m)})\big)\right].$ (21)
@@
$\mathcal{L}_{\text{total}}=\lambda_{\text{match}}\mathcal{L}_{\text{match}}+\lambda_{\text{RL}}\mathcal{L}_{\text{RL}},$ (22)
where $\lambda_{\text{match}}$ and $\lambda_{\text{RL}}$ are weighting coefficients balancing the contributions of trajectory matching and policy optimization.
Our reinforcement-augmented training framework addresses key limitations of diffusion-based planners in end-to-end autonomous driving. Through reward-guided GRPO optimization, we impose safety, diversity, and trajectory consistency constraints that are not captured by conventional MSE-based training. This hybrid supervision strategy bridges the gap between behavior cloning and policy-level optimization, leading to more stable, safe, and deployable planning performance.
@@ 5.2 Evaluation Metrics
Overall, a higher Diversity value indicates greater dispersion among the predicted trajectories, reflecting a richer coverage of intent modes such as lane changes, braking, and yielding. When Diversity approaches 1, the predicted modes are highly diverse and well-separated in the trajectory space. Conversely, a value near 0 suggests that the predicted trajectories are highly similar, indicating a lack of behavioral diversity. Notably, the Diversity Metric is only applicable to multi-mode trajectory E2E-AD methods.
@@ 5.3 Implementation Details
[…]
@@
WoTE* [68] ST IL ICCV 2025 - - 61.71 31.36 - - - - - - - -
@@
Figure 6: Visualization of Planning Loss and Reinforcement Learning Loss of DIVER on the nuScenes [54] dataset, showing stable convergence without oscillation or divergence.
Figure 8: Visualization results of DIVER compared with SparseDrive [6] on the Bench2Drive [53] dataset. It further confirms DIVER’s powerful capability in generating diverse trajectories.
TABLE II: Comparison on planning-oriented $\operatorname{NAVSIM}$ navtest split with $\operatorname{\textbf{Closed-Loop}}$ metrics. ‘mmt’ refers multi-mode trajectory variant of $\operatorname{TransFuser}$ and denotes the re-implementation.
@@
[…]
@@
TABLE IX: Ablation Study on Policy-Aware Diffusion Generator on NAVSIM dataset. ‘Condition’ denotes map and agent conditioning in the diffusion model, while ‘Reference GTs’ refers to reference trajectory guidance. ‘FPS’ is measured on an NVIDIA 4090 GPU.
@@
DiffusionDrive 98.2 96.2 94.7 100 82.2 88.1 45
$\operatorname{TransFuser}_{\operatorname{mmt}}$$^{*}$ 96.2 95.4 90.7 100 80.7 85.1 56
+ Condition 97.2 96.3 90.8 100 81.2 86.3 44
+ Reference GTs 97.6 95.7 91.5 100 81.6 86.4 50
$\operatorname{DIVER}$ 98.5 96.5 94.9 100 82.6 88.3 41
@@
TABLE XI: Ablation Study on the TrajPooler on NAVSIM dataset. FullImg denotes full image features without projection and NoImg denotes no image features (trajectory encoding only).
[…]
DIVER+NoImg 96.9 95.6 94.1 100 81.5 87.7
DIVER+FullImg 97.3 95.9 94.5 100 81.9 87.9
DIVER+TrajPooler 98.5 96.5 94.9 100 82.6 88.3
TABLE XII: Ablation Study on Loss Function Design on NuScenes dataset.
@@
TABLE XIII: Ablation Study on Loss Function Design on NAVSIM dataset.
@@
TABLE XIV: Ablation Study on different reward on Bench2Drive dataset. We use SparseDrive as the baseline.
$\operatorname{Reward}$ $\operatorname{Open-loop\ Metric}$ $\operatorname{Closed-loop\ Metric}$
$r_{\text{div}}$ $r_{\text{safe}}$ $r_{\text{TC}}$ $r_{\text{LK}}$ $\operatorname{Avg.\textit{Div.}^{(t)}}\uparrow$ $\operatorname{DS}\uparrow$ $\operatorname{SR\ (\%)}\uparrow$ $\operatorname{Effi}\uparrow$ $\operatorname{Comf}\uparrow$
0.21 44.54 16.71 170.21 48.63
$\checkmark$ $\checkmark$ 0.35 49.21 21.56 177.00 54.72
$\checkmark$ $\checkmark$ $\checkmark$ 0.33 48.71 21.40 176.81 56.10
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.32 48.89 21.37 176.34 56.23
TABLE XV: Ablation Study on different reward on NuScenes dataset. We use SparseDrive as the baseline.
$\operatorname{Reward}$ $\operatorname{\textit{Div.}^{(t)}}\uparrow$ $\operatorname{Col.\ Rate\ (\%)}\downarrow$
$r_{\text{div}}$ $r_{\text{safe}}$ $r_{\text{TC}}$ $r_{\text{LK}}$ 1s 2s 3s $\operatorname{Avg.}$ 1s 2s 3s $\operatorname{Avg.}$
0.05 0.11 0.23 0.13 0.01 0.05 0.18 0.08
$\checkmark$ $\checkmark$ 0.10 0.19 0.34 0.21 0.01 0.05 0.15 0.07
$\checkmark$ $\checkmark$ $\checkmark$ 0.10 0.18 0.35 0.20 0.01 0.05 0.16 0.07
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.09 0.17 0.35 0.20 0.01 0.05 0.16 0.07
TABLE XVI: Ablation Study on different reward on NAVSIM dataset. We use $\operatorname{TransFuser}_{\operatorname{mmt}}$$^{*}$ as the baseline.
[…]
96.2 95.4 90.7 100 80.7 85.1
$\checkmark$ $\checkmark$ 98.5 96.5 94.9 100 82.6 88.3
$\checkmark$ $\checkmark$ $\checkmark$ 98.9 96.6 94.7 100 82.3 88.6
@@ 5.4 Main Results
To comprehensively evaluate the multi-mode trajectory generation capability of our proposed DIVER, we conduct closed-loop experiments on two representative E2E-AD benchmarks: Bench2Drive and NAVSIM, as shown in Tables I and II. In Table III, we further present open-loop results on the NuScenes dataset. Additionally, Table I includes supplementary open-loop evaluation results on Bench2Drive. It is important to note that we do not adopt the conventional L2 distance in open-loop evaluation. This is because L2 distance merely reflects the proximity between a predicted trajectory and a single GT trajectory, failing to capture the diversity and spread of multi-mode trajectory predictions. Relying solely on this metric compromises the core objective of multi-mode trajectory generation—namely, to cover a wide range of plausible future intents.
@@ 5.4.1 Close-Loop Results
NAVSIM (Close-Loop). Table II reports the closed-loop evaluation results on the NAVSIM dataset, where we follow the official planning-oriented protocol [59]. The metrics include No at-fault Collisions (NC), Drivable Area Compliance (DAC), Time-to-Collision (TTC), Comfort (Comf.), Ego Progress (EP), and PDM score (PDMS), which jointly assess the safety, smoothness, and efficiency of long-horizon driving behaviors. Our proposed DIVER achieves the best overall performance across nearly all metrics, attaining a PDMS score of 88.3, which surpasses the strong diffusion-based baseline $\operatorname{DiffusionDrive}$ (88.1) and the multi-mode $\operatorname{TransFuser}_{\text{mmt}}$ (85.1). DIVER also records the highest NC (98.5) and DAC (96.5) scores, indicating superior collision avoidance and lane compliance. Moreover, it maintains full ride comfort (100) while achieving competitive ego progress (82.6), demonstrating a well-balanced capability between safety and efficiency. Compared with Hydra-MDP [11], which distills PDMS metrics into a static policy in a single-step process, DIVER dynamically incorporates PDMS-guided reinforcement within the diffusion-based trajectory generation. These results highlight DIVER’s ability to balance safety, comfort, and goal-directed planning, demonstrating its effectiveness in realistic driving simulations. The improvements across all aspects suggest that our method not only generates diverse multi-mode trajectories but also ensures their practical executability and robustness in closed-loop scenarios.
@@ 5.6.1 Ablation Study on PADG
Impact of the Number of Reference GTs. As shown in Figure 5, increasing the number of Reference GTs leads to consistent gains in Driving Score, Success Rate, and Merging. Performance stabilizes around 6–7 GTs, indicating that moderate reference diversity effectively enhances planning robustness and behavioral richness.
Impact of the TrajPooler module. TrajPooler is designed to focus on image regions that are spatially correlated with the predicted trajectories rather than processing the entire image. As shown in Table XI, to evaluate the effectiveness of the proposed TrajPooler module, we conduct ablation experiments comparing three configurations: full image features without projection, no image features (trajectory encoding only), and the TrajPooler module. These results indicate that TrajPooler effectively aggregates perception features around the predicted trajectory, leading to a 1.2% increase in NC and a 0.4% increase in PDMS compared with using full image features, thereby enabling the model to focus on spatially relevant cues while avoiding redundant global contexts. The results demonstrate that TrajPooler not only reduces computational cost but also enhances the model’s ability to focus on trajectory-relevant regions, leading to improved representation efficiency.
@@ 5.6.2 Ablation Study on Loss Function Design
Impact of Different Loss Function. As shown in Tables XII and XIII, loss function design has a significant impact on model performance. On the NuScenes dataset, the standard L1 loss fails to exploit the potential of diffusion and reference GTs, yielding limited diversity (Avg. Div.↑ = 0.14). Even when combined with RL (PPO), improvements are marginal. In contrast, replacing L1 with the proposed matching loss $\mathcal{L}_{\text{match}}$ enables the model to benefit from trajectory diversity, yielding higher $\operatorname{\textit{Div.}^{(t)}}$ (up to 0.21). Further integrating RL with diversity and safety rewards enhances performance, especially using GRPO, which outperforms PPO in both diversity and planning quality. On NAVSIM, $\mathcal{L}_{\text{match}}+\mathcal{L}_{\text{RL(GRPO)}}$ achieves the best PDMS (88.3), confirming that this design effectively balances realism, diversity, and safety.
As shown in Figure 6, we visualize the dynamics of the planning loss and the reinforcement learning loss during training. The results show that both losses converge smoothly without oscillation or divergence, indicating stable optimization. By jointly leveraging reinforcement learning and imitation learning, our model mitigates the mode collapse commonly observed in pure imitation learning approaches and alleviates the gradient explosion issue that often occurs in standalone reinforcement learning models. Impact of Different Reward Mechanisms. As shown in Tables XIV, XV, and XVI, incorporating $r_{\text{div}}$ and $r_{\text{safe}}$ leads to notable quantitative improvements across all datasets. On Bench2Drive, the average diversity $\operatorname{Div.}^{(t)}$ increases from 0.21 to 0.35 (+66.7%), and the success rate rises from 16.71% to 21.56% (+29.1%). On NuScenes, the average trajectory diversity improves from 0.13 to 0.21 (+61.5%) while maintaining a low collision rate (0.08→0.07). On NAVSIM, combining $r_{\text{div}}$, $r_{\text{safe}}$, and PDMS guidance yields the best performance, with NC improving from 96.2 to 98.9 (+2.7%) and DAC from 95.4 to 96.6 (+1.3%). These quantitative results confirm that multi-term reward design significantly enhances both trajectory diversity and driving safety across benchmarks.